OpenStudy (anonymous):

How would I find the kernel of the linear transformation and express the answer as the span of a set of vectors? For A=[2 1 -1; 1 -2 1; 1 -7 4; 3 4 -3]

OpenStudy (anonymous):

I reduced it to row echelon form and got two independent variables [1 -7 4; 0 1 -3/5; 0 0 0; 0 0 0]. Then I got stuck from there. I don't know how to set the equations up for this matrix.

OpenStudy (anonymous):

Oh my bad the ref I got was [1 0 -1/5; 0 1 -3/5; 0 0 0; 0 0 0]

OpenStudy (anonymous):

when you row reduce the the matrix, you identify the pivot columns and the free variable columns. Then you write the pivot variables in terms of the free variables, and factor out the free variables to get the vectors that span the Null Space (or Kernel).

OpenStudy (anonymous):

Why is it sometimes the last column is not a variable column?

OpenStudy (anonymous):

I get confused when that applies

OpenStudy (anonymous):

you mean, not a free variable column? if the last column isnt free, and they are all pivot columns, that means those 3 columns are linearly independent, and the Null Space (Kernel) will just be the 0 vector.

OpenStudy (anonymous):

Umm I don't know if we're talking about the same thing like let's say it's [1 0 5; 0 1 4; 0 0 0] sometimes the answer becomes x_1 = 5 and x_2 =4?

OpenStudy (anonymous):

And why is it for the answer there's only 3 variables? If there's two rows of zeroes aren't there 2 free variables?

OpenStudy (anonymous):

in that example, when we turn that matrix back into a system of a equations it would look like this: \[x_1+5x_3= 0 \] \[x_2+4x_3= 0\] \[0 = 0\] Its a 3x3 matrix, so thats 3 equations in 3 unknowns. The pivot variables are x1 and x2, the free variable is x3.

OpenStudy (anonymous):

The reason for the zeros on the right side of the equation comes from the fact we are looking for the kernel, which is finding the solutions to Ax = 0

OpenStudy (anonymous):

So to complete the problem, we solve for the pivot variables: \[x_1+5x_3 = 0 \Rightarrow x_1 = -5x_3\] \[x_2+4x_3=0\Rightarrow x_2 = -4x_3\] x3 is free. so the vector looks like: -5x3 -4x3 x3 and when you factor out the x3 you get the vector: -5 -4 1

OpenStudy (anonymous):

Oh okay that make sense thank you, I forgot about the Ax=0. But last question so even if you have two rows of zeroes that would count as 1 free variable.

OpenStudy (anonymous):

When would you have two sets of basis for the kernel T?

OpenStudy (anonymous):

yes, that would still just be one free variable. it doesnt matter how many zero rows you have. I think you are asking when would there be two vectors that span the Kernel of a matrix? That would be the case if there were 2 free variables. The number of vectors you get for your basis of the Kernel = number of free variables.

OpenStudy (anonymous):

Oh okay thank you! That makes a lot more sense

OpenStudy (anonymous):

Actually sorry here's what I'm confused about like for this example in my book it says A=[1 -8 -7 -4; 2 -3 -1 5; 3 2 5 14] and when reduced it becomes [1 0 1 4 -8; 0 1 1 1 -2; 0 0 0 0]

OpenStudy (anonymous):

x_1 = -8 s - 4t x_2 = -2 -s -t x_3 = s x_4 =t

OpenStudy (anonymous):

Any ideas why the last column isn't a free variable column? Oh and this question isn't asking for a the kernel for whether b is the column space of A

OpenStudy (anonymous):

So is it only if it's asking for the kernel the last column is also a free variable column?

OpenStudy (anonymous):

i dont like your books notation =/ anywhos, When we are looking for the Column space of a matrix, we are only looking for the pivot columns. So we row reduce a matrix, see which columns have pivots, then look back at the original problem to see which columns those were, and thats a basis for the column space. Those columns span the columns space. When looking for the Kernel, thats when we are interested in the free variables instead. Im looking at the problem you posted, writing out a solution on paper.

OpenStudy (anonymous):

Oh my bad I get what they did

OpenStudy (anonymous):

They gave a b vector too b=[8 -10 -28]

OpenStudy (anonymous):

b=[8; -10; -28]

OpenStudy (anonymous):

Oh your explanation actually makes a lot more sense than what it states in the textbook!! Thanks a lot man, you really know your linear algebra material haha

OpenStudy (anonymous):

Oh I see so it only works if all solutions exists

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (anonymous):

no problem :)